OpenStudy (anonymous):

how would you solve 1/3+1/5=1/x

OpenStudy (lgbasallote):

hmm correct me if im wrong but.. \[\large \frac{1}{x} = \frac 13 + \frac 15 \implies x = 3 + 5\] right?

OpenStudy (mimi_x3):

Is it: \[\large \frac{1}{3} +\frac{1}{5} =\frac{1}{x} ?\]

OpenStudy (mimi_x3):

something is wrong igba

OpenStudy (anonymous):

\[\frac{1}{3}+\frac{1}{5}=\frac{1}{x} \]\[\frac{8}{15}=\frac{1}{x}\]\[x=\frac{15}{8} \]

OpenStudy (lgbasallote):

hmm yep something was wrong lol

OpenStudy (anonymous):

if the q is like that then the answer shouldn't be an integer 1/3 + 1/5 = 5/15 +3/15 = 8/15 to make the numerator 1, divide it by 8 15/8 = 1.875

OpenStudy (mimi_x3):

\[\frac{1*5}{3*5} +\frac{1*3}{5*3} =\frac{1}{x} => \frac{5+3}{15} = \frac{1}{x} => \frac{8}{15} =\frac{1}{x} => 8x= 15 => x=\frac{15}{8} \]

OpenStudy (anonymous):

thanks guys :)